If f(x)=2x^2-3x-6, find the value of f(-3)+f(2). (I'm completely stumpted on this one :( )
Answer by ewatrrr(24785)


Hi
f(x)=2x^2-3x-6
 f(2) just mean to substitute 2 for x in the equation.
 f(-3) + f(2) = 2*9 + 9 - 6   +    2*4 -6 - 6 = 17

Wish You the Best in your Studies.